# Improving the Validity of Automatically Generated Feedback via Reinforcement Learning

### v1

@@ 5 Conclusions and Future Work
In this work, we proposed a framework for automated feedback generation and evaluation via LLMs for students’ incorrect answers in math multiple-choice questions. Our framework accounts for both the mathematical correctness of the feedback and its alignment with good pedagogical practices. We show that using a data augmentation and preference optimization approach, we can generate high-quality feedback using Llama 2 7B, a small and open-source LLM. We also show that GPT-4 can evaluate feedback rather accurately using a rubric and that its annotations are helpful for training the feedback generation method. There are many avenues for future work. First, we can apply our framework to other RL algorithms such as PPO, or non-RL approaches such as overgenerate-and-rank. Second, we can evaluate our final feedback generation task via a large-scale human evaluation or classroom study, which would alleviate concerns on GPT-4’s annotations being biased. Finally, we can test our framework’s generalizability by applying it to domains other than math, such as programming or language learning, or other scenarios such as hint generation or student-instructor conversations.

### v2

@@ 5 Conclusions and Future Work
In this work, we proposed a framework for automated feedback generation and evaluation via LLMs for students’ incorrect answers in math multiple-choice questions. Our framework accounts for both the mathematical correctness of the feedback and its alignment with good pedagogical practices. We show that using a data augmentation and preference optimization approach, we can generate high-quality feedback using Llama 2 7B, a small and open-source LLM. We also show that GPT-4 can evaluate feedback rather accurately using a rubric and that its annotations are helpful for training the feedback generation method. There are many avenues for future work. First, we can apply our framework to other RL algorithms such as PPO, or non-RL approaches such as overgenerate-and-rank. Second, we can evaluate our final feedback generation task via a large-scale human evaluation or classroom study, which would alleviate concerns on GPT-4’s annotations being biased. Third, we can test our framework’s generalizability by applying it to other domains such as programming or language learning, or other scenarios such as hint generation or student-instructor conversations. Finally, we can consider tailoring feedback to each student according to their knowledge levels [19], especially for open-ended questions, since student errors can likely be detected from these responses [21, 36, 37].
Acknowledgments
The authors would like to thank Jaewook Lee and Hunter McNichols for assisting in the presentation of this work. The authors also thank Schmidt Futures and the NSF (under grants IIS-2118706 and IIS-2237676) for partially supporting this work.